OpenStudy (anonymous):

the supplement of an angle is 10 more than twice its complement. Find the angle.

OpenStudy (anonymous):

10 is my answer. Am I right?

OpenStudy (raden):

let there is an angle x. the supplement of x is (180 - x) and its complement 90 - x. according information above, we have an equation : 180 - x = 2(90 - x) + 10 solve for x

OpenStudy (anonymous):

so.. 2(90-x)+10 -x= -2x+10 x= 10 am I right, @RadEn ?

OpenStudy (raden):

oh, yes you are right :)

OpenStudy (anonymous):

Thank you. :)

OpenStudy (raden):

welcome